# Capturing Human Motion from Monocular Images in World Space with Weak-supervised Calibration

### v4

@@ IV-D Evaluation Results
Camera Coordinate In our study, we choose three challenging datasets containing distorted images to showcase the superiority of weak-supervised camera calibration under complex conditions. As depicted in Tab. II, our method surpasses the performance of most existing methods on the challenging synthetic dataset AGORA [42]. Each of our models demonstrates unique strengths. The pure regression model achieves higher accuracy in joint localization, and the hybrid regression module excels in capturing mesh details. Compared to baseline [13], our improvement is quite substantial. Note the baseline’s backbone, training datasets, and evaluation settings are the same as Ours. As depicted in Fig. 6, our W-HMR outperforms SPEC [3] in handling occlusions, rare body rotations, and body intersections. Although the joint accuracy does not exhibit significant improvement on the indoor dataset HuMMan [44], we achieve the closest performance compared to the latest method Zolly in PVE, even surpassing the model-free method FastMETRO [54]. The most remarkable performance is observed on the real distorted dataset SPEC-MTP. We significantly elevate the SOTA results on SPEC-MTP to new heights.
TABLE III: Comparison with SOTA methods on SPEC-MTP [3]. The first six models in the table are all methods that output results in the camera coordinate, so there are two results. The ”a” in ”a/b” indicates the result calculated directly using the output in the camera coordinate. The ”b” is the result calculated after camera rotation predicted by CamCalib has calibrated the output. The latter two models can directly get the results in the world coordinate without calibration, so there is only one result.
@@
GraphCMR [53] 175.1/166.1 94.3 205.5/197.3
SPIN [9] 143.8/143.6 79.1 165.2/165.3
PartialHumans [59] 158.9/157.6 98.7 190.1/188.9
I2L-MeshNet [30] 167.2/167.0 99.2 199.0/198.1
HMR [1] 142.5/128.8 71.8 164.6/150.7
PyMAF [13] 148.8/134.2 66.7 166.7/158.7
@@
World Coordinate We utilize SPEC-MTP [3] as the primary evaluation dataset. Due to its only test set, specific fine-tuning is not feasible. It adopts a shooting method different from traditional datasets, resulting in unique rotation and severe distortion. Therefore, it poses a significant challenge and is suitable for demonstrating OrientCorrect’s strength. We only evaluate the W-HMR with a pure parameter regression module to ensure fairness. W-HMR easily outperforms existing methods, including SPEC. It is worth noting that the superior performance of W-HMR in the world coordinates is not entirely dependent on higher reconstruction accuracy. Perhaps you noticed two Ours-Ps in Tab. III. This is because W-HMR can generate results in two coordinates simultaneously. By comparing these two sets of results, we can conclude that OrientCorrect is significantly superior to the traditional method of directly using predicted camera rotation for correction. Upon comparing our results calibrated by CamCalib with those corrected by OrientCorrect, it is evident that our module is effectively functioning. As illustrated in Fig. 1, the reconstructed human body pose appears implausible in world space when the camera coordinate is not aligned with the world coordinate. However, our OrientCorrect corrects the orientation of the human body and maintains its reasonableness in the world coordinate. Our model can accurately perceive and correct both slight and significant camera rotation. Quantitative and qualitative results demonstrate that our method is simpler and more efficient than SPEC, which will be further substantiated in subsequent ablation studies.
Figure 7: Qualitative results on HuMMan [44]. ”Input” refers to input images of the model. ”Overlay” means we project the recovered mesh back onto the original images. ”Cam” and ”World” show the output in camera and world coordinates, respectively.
TABLE IV: Comparison with SOTA Methods on 3DPW [39]. The column ”w/ 3DPW” indicates whether the model has been trained using the 3DPW dataset. We bold both the best results with and without 3DPW.
@@
HMR [1] No 130.0 76.7 -
GraphCMR [53] No - 70.2 -
SPIN [9] No 96.9 59.2 116.4
SPEC [3] Yes 96.4 52.7 -
I2L-MeshNet [30] No 93.2 57.7 110.1
Graphomer [12] Yes 74.7 45.6 87.7
FastMETRO [54] Yes 73.5 44.6 84.1
PyMAF [13] Yes 92.8 58.9 110.1
VIBE [10] No 93.5 56.5 113.4
VIBE [10] Yes 82.9 51.9 99.1
SATF [60] No 81.2 48.7 96.0
STAF [60] Yes 80.6 48.0 95.3
PARE [11] No 84.1 49.3 99.4
PARE [11] Yes 79.1 46.4 94.2
HybrIK [56] No 80.0 48.8 94.5
HybrIK [56] Yes 74.1 45.0 86.5
Zolly [2] No 76.2 47.9 89.8
Zolly [2] Yes 65.0 39.8 76.3
@@
TABLE V: Comparison with SOTA Methods on controllable datasets including Human 3.6M [38] and MPI-INF-3D [40].
Model H36M MPI-INF-3D
[…]
HMR [1] 88.0 56.8 124.2 89.8
GraphCMR [53] - 50.1 - -
SPIN [9] 62.5 41.1 105.2 67.5
PyMAF [13] 57.7 40.5 - -
FastMETRO [54] 53.9 37.3 - -
HMMR [48] - 56.9 - -
VIBE [10] 78.0 53.3 96.6 64.6
MEVA [61] 76.0 53.2 96.4 65.4
HybrIK [56] 54.4 34.5 - -
CLIFF [4] 47.1 32.7 - -
MPS-Net [49] 69.4 47.4 96.7 62.8
STAF [60] 70.4 44.5 92.4 58.8
Zolly [2] 49.4 32.3 - -
Ours 52.7 34.0 88.3 62.0
Ours 51.0 29.8 83.2 59.1
Ours-P 45.5 30.2 83.3 59.8
Traditional Benchmarks Following previous methods [13, 49], we conduct evaluation on 3DPW [39], H36M [38], and MPII-INF-3D [40]. Specifically, we evaluate the performance of models trained with and without the 3DPW train set on the 3DPW test set. As shown in Tab. IV, we achieve remarkable improvements on this challenging in-the-wild dataset, surpassing previous works easily. Even compared to the SOTA method Zolly, W-HMR still exhibits significant advancements when trained without the 3DPW train set, which shows W-HMR’s robustness. In Tab. V, W-HMR’s performance is relatively less impressive in the case of controllable datasets, although it remains competitive with these SOTA methods. Controllable datasets refer to datasets obtained under laboratory conditions, typically characterized by limited variations in actors and scenes, which are prone to overfitting. HuMMan in Tab. II is an example of such datasets. So, it becomes difficult for our advantages to be fully demonstrated.
@@
Figure 9: Qualitative comparison with SPEC on SPEC-MTP [3]. Here are some challenging examples. Although imperfect, W-HMR shows better mesh-image alignment and higher orientation accuracy in world space.
Camera Parameter There is no such evaluation for inferred camera parameters because W-HMR cannot and does not need to predict true intrinsic and extrinsic parameters. First, focal length prediction is an ill-posed problem and lacks much training data. We choose to leverage mesh-image alignment information to predict pseudo focal lengths. This pseudo focal length may differ from the true focal length but can accurately project the recovered mesh to the human area. After all, the purpose of predicting the focal length is to handle distorted images so that the calculation of 2D joint loss is more accurate, thereby improving model performance. In summary, inferred camera parameters cannot be evaluated quantitatively because they may differ significantly from ground truth. Of course, it can be implicated by improved recovery accuracy and aligned-well qualitative results. Moreover, the module used to predict the camera rotation matrix in W-HMR comes from CamClib in SPEC [3]. Its accuracy has been evaluated in SPEC. The novelty in our work is how we better use this information. With the same CamClib, W-HMR still significantly outperforms SPEC, which is a solid evidence to prove the effectiveness of our method.
IV-E Ablation Study
@@
TABLE VI: The ablation study on the core block of the W-HMR. The experiments below utilize ResNet50 [36] as the backbone. In this table, ”V” refers to the vertex regressor. ”F” indicates whether the focal length predictor and FULL2 2D supervision are implemented. ”G” signifies the addition of global information during input.
@@
Camera Calibration We select PyMAF[13] as our baseline and consistently use ResNet50 as the backbone throughout our ablation study, with 431 sparse vertices as the downsampling points. The training and evaluation settings are all kept the same. From the Tab. VI, it is evident that the most significant improvements stem from adding the vertex regressor and global information. Including the vertex regressor enhances the model’s ability to reconstruct body mesh details. Moreover, global information contributes independently and aids in camera calibration. A comparison between ”V” and ”V+F” reveals that solely adding the focal length predictor only has a noticeable impact on vertex accuracy. However, comparing ”V+G” and ”V+F+G” demonstrates that the presence of global information leads to more substantial improvements when combined with the focal length predictor. This finding further supports our idea that global information is crucial for recovering body orientation, predicting focal lengths, and correcting distortions. In summary, with the camera calibration assistance, W-HMR achieves a decent reconstruction accuracy, which provides a solid foundation for human body recovery in world space.
TABLE VII: The ablation study on body orientation correction. ”spec-like” refers to using a similar method like SPEC to predict global pose directly. The numbers in brackets are the evaluation results on SPEC-MTP. This is the reason why we do not choose the spec-like method. Its impressive results are based on overfitting and sacrificing generalization.
@@
GraphCMR [53] 181.7/181.5 86.6 219.8/218.3
SPIN [9] 165.8/161.4 79.5 194.1/188.0
PartialHumans [59] 169.3/174.1 88.2 207.6/210.4
I2L-MeshNet [30] 169.8/163.3 82.0 203.2/195.9
HMR [1] 128.7/96.4 55.9 144.2/111.8
PyMAF [13] 126.8/88.0 48.7 136.7/118.0
@@
Orientation Correction We select two methods for evaluation. The first method is similar to SPEC. SPEC concatenates features and the camera parameters obtained from CamCalib, then inputs them into a regressor to directly recover the human body in the world coordinate. As demonstrated in Tab. VII, the spec-like method significantly outperforms previous methods. However, we attribute this performance to overfitting. The model relies on massive parameters to memorize the global body orientation in the training set of SPEC-SYN, subsequently achieving excellent results on the test set. This is supported by the inferior performance of spec-like methods on SPEC-MTP. This indicates that such a method is prone to overfitting, leading to the misconception that the model can capture environmental information. Worse, this design destroys the model’s original learned knowledge, making the overall reconstruction less precise. In contrast, our method offers a more straightforward structure, requires less training data, and exhibits greater robustness. Furthermore, other models based on the camera coordinate can easily adopt our approach to achieve global pose recovery.
TABLE VIII: Ablation study results on SPEC-MTP. The weights are from SPIN and are not fine-tuned. indicates using orientation output by OrientCorrect of pre-trained W-HMR. indicates HMR with an adaptive OrientCorrect module.
@@
To verify the portability of OrientCorrect, we do related experiments, and the results are displayed in Tab. VIII. To make the experiments more convincing, we choose HMR, a widely adopted model, as the baseline. We design two porting methods. One is to directly use the pre-trained OrientCorrect from W-HMR to predict the body rotation in the world space and then combine it with the pose and shape output from HMR to generate the final human motion. It can be seen that it brings nearly 14% improvement. From Tab. III, it can be seen that using the camera rotations output from SPEC’s CamCalib to correct brings only about 10% boost.
Moreover, the most impressive is the second porting method. We utilize the features and body orientation output by HMR to replace the original input part of OrientCorrect. In other words, we train an OrientCorrect adapted to HMR. The effect is pronounced, and the reconstruction accuracy is again improved by one grade. To avoid misunderstandings, you may think OrientCorrect is not as good as CamCalib or SPEC because the results in Tab. VIII are not as good as those of HMR and SPEC in Tab. III. This is because we are based on different baselines. We use the pre-trained weights from SPIN directly and do not do any tuning, so the final results should be worse.
Design Details To find the optimal design for our model, we conduct further ablation experiments, as shown in Tab. IX. Firstly, we test different downsampling methods. Using markers for downsampling leads to better learning of human body features. Here, ”markers” refer to mocap markers. The improved performance might be because mocap markers are a better proxy for inferring human body information [62]. However, since using mocap markers conflicts with the vertex regressor, we only employ markers in the pure parameter regression.
TABLE IX: Ablation experiments regarding the design details of modules. The first comparison involves vertex selection for downsampling. The second comparison examines the impact of different information inputs. The third comparison evaluates different designs of the vert graphomer. The fourth comparison evaluates different combinations of regressors. Within each comparison, all settings remain consistent except for the variations in module design.
@@
We also compare our 5D global information with the 3D bounding box information proposed by CLIFF [4]. It is clear that, apart from the relative position information provided by the bounding box, the original image size also contains essential information for our task.
@@
Lastly, we test various combinations of the hybrid regression module. Both approaches show advantages, but it is apparent that the combination of ”2 param$+$1 vert” has clear benefits regarding model size and inference efficiency. Therefore, we ultimately select this combination as our final regression module.
@@
V Discussion
Difference It is not the first time that predicting the distance between a person and a camera has been proposed. To understand our innovation, we take [63] as an example to discuss the differences between our work. Different Methods: The neural network in our work directly predicts the person’s distance from the camera. But in [63], they first calculate values $k$ based on the sizes of the bounding box and image. Then, they predict a correction factor $\gamma$ through a neural network and get the absolute depth value using $k$ and $\gamma$. Different Theories: Our work mainly relies on the degree of distortion of humans in the images. The model tends to predict shorter distances when humans are more distorted. But [63] mainly relies on the human size in images. A more miniature human means this human stands farther away from the camera. Different Purposes: We use predicted distance to calculate focal lengths. Our ultimate goal is to achieve a perspective camera model so that we can make full use of distorted images. As for [63], they want to recover accurate relative positions for multi-person pose estimation.

### v6

@@ 4.3 Evaluation Results
Camera Coordinate In our study, we choose three challenging datasets containing distorted images to showcase the superiority of weak-supervised camera calibration under complex conditions. As depicted in Tab. 1, our method surpasses the performance of most existing methods on the challenging synthetic dataset AGORA (Patel et al., 2021). Each of our models demonstrates unique strengths. The pure regression model achieves higher accuracy in joint localization, and the hybrid regression module excels in capturing mesh details. Compared to baseline (Zhang et al., 2021a), our improvement is quite substantial. Note the baseline’s backbone, training datasets, and evaluation settings are the same as Ours. Although the joint accuracy does not exhibit significant improvement on the indoor dataset HuMMan (Cai et al., 2022), we achieve the closest performance compared to the latest method Zolly in PVE, even surpassing the model-free method FastMETRO (Cho et al., 2022). The most remarkable performance is observed on the real distorted dataset SPEC-MTP. We significantly elevate the SOTA results on SPEC-MTP to new heights.
Table 2: Comparison with SOTA methods on SPEC-MTP (Kocabas et al., 2021b). The first six models in the table are all methods that output results in the camera coordinate, so there are two results. The “a” in “a/b” indicates the result calculated directly using the output in the camera coordinate. The “b” is the result calculated after camera rotation predicted by CamCalib has calibrated the output. The latter two models can directly get the results in the world coordinate without calibration, so there is only one result.
@@
GraphCMR 175.1/166.1 94.3 205.5/197.3
SPIN 143.8/143.6 79.1 165.2/165.3
PartialHumans 158.9/157.6 98.7 190.1/188.9
I2L-MeshNet 167.2/167.0 99.2 199.0/198.1
HMR 142.5/128.8 71.8 164.6/150.7
PyMAF 148.8/134.2 66.7 166.7/158.7
@@
World Coordinate We utilize SPEC-MTP (Kocabas et al., 2021b) as the primary evaluation dataset. Due to its only test set, specific fine-tuning is not feasible. It adopts a shooting method different from traditional datasets, resulting in unique rotation and severe distortion. Therefore, it poses a significant challenge and is suitable for demonstrating OrientCorrect’s strength. We only evaluate the W-HMR with a pure parameter regression module to ensure fairness. W-HMR easily outperforms existing methods, including SPEC. It is worth noting that the superior performance of W-HMR in the world coordinates is not entirely dependent on higher reconstruction accuracy. Perhaps you noticed two Ours-Ps in Tab. 2. This is because W-HMR can generate results in two coordinates simultaneously. By comparing these two sets of results, we can conclude that OrientCorrect is significantly superior to the traditional method of directly using predicted camera rotation for correction. Upon comparing our results calibrated by CamCalib with those corrected by OrientCorrect, it is evident that our module is effectively functioning. As illustrated in Fig. 1, the reconstructed human body pose appears implausible in world space when the camera coordinate is not aligned with the world coordinate. However, our OrientCorrect corrects the orientation of the human body and maintains its reasonableness in the world coordinate. Our model can accurately perceive and correct both slight and significant camera rotation. Quantitative and qualitative results demonstrate that our method is simpler and more efficient than SPEC, which will be further substantiated in subsequent ablation studies.
Table 3: Comparison with SOTA Methods on controllable datasets including Human 3.6M (Ionescu et al., 2013) and MPI-INF-3D (Mehta et al., 2017a).
Model H36M MPI-INF-3D
[…]
HMR 88.0 56.8 124.2 89.8
GraphCMR - 50.1 - -
SPIN 62.5 41.1 105.2 67.5
PyMAF 57.7 40.5 - -
FastMETRO 53.9 37.3 - -
HMMR - 56.9 - -
VIBE 78.0 53.3 96.6 64.6
MEVA 76.0 53.2 96.4 65.4
HybrIK 54.4 34.5 - -
CLIFF 47.1 32.7 - -
MPS-Net 69.4 47.4 96.7 62.8
STAF 70.4 44.5 92.4 58.8
Zolly 49.4 32.3 - -
Ours 52.7 34.0 88.3 62.0
Ours 51.0 29.8 83.2 59.1
Ours-P 45.5 30.2 83.3 59.8
Traditional Benchmarks Following previous methods (Zhang et al., 2021a; Wei et al., 2022), we conduct evaluation on 3DPW (Von Marcard et al., 2018), H36M (Ionescu et al., 2013), and MPII-INF-3D (Mehta et al., 2017a). Specifically, we evaluate the performance of models trained with and without the 3DPW train set on the 3DPW test set. As shown in Tab. 4, we achieve remarkable improvements on this challenging in-the-wild dataset, surpassing previous wor-ks easily. Even compared to the SOTA method Zolly, W-HMR still exhibits significant advancements when trained without the 3DPW train set, which shows W-HMR’s robustness. In Tab. 3, W-HMR’s performance is relatively less impressive in the case of controllable datasets, although it remains competitive with these SOTA methods. Controllable datasets refer to datasets obtained under laboratory conditions, typically characterized by limited variations in actors and scenes, which are prone to overfitting. HuMMan in Tab. 1 is an example of such datasets. So, it becomes difficult for our advantages to be fully demonstrated.
Table 4: Comparison with SOTA Methods on 3DPW (Von Marcard et al., 2018). The column “w/ 3DPW” indicates whether the model has been trained using the 3DPW dataset. We bold both the best results with and without 3DPW.
@@
HMR No 130.0 76.7 -
GraphCMR No - 70.2 -
SPIN No 96.9 59.2 116.4
SPEC Yes 96.4 52.7 -
I2L-MeshNet No 93.2 57.7 110.1
Graphomer Yes 74.7 45.6 87.7
FastMETRO Yes 73.5 44.6 84.1
PyMAF Yes 92.8 58.9 110.1
VIBE No 93.5 56.5 113.4
VIBE Yes 82.9 51.9 99.1
SATF No 81.2 48.7 96.0
STAF Yes 80.6 48.0 95.3
PARE No 84.1 49.3 99.4
PARE Yes 79.1 46.4 94.2
HybrIK No 80.0 48.8 94.5
HybrIK Yes 74.1 45.0 86.5
Zolly No 76.2 47.9 89.8
Zolly Yes 65.0 39.8 76.3
@@
Figure 7: Qualitative results on HuMMan (Cai et al., 2022). “Input” refers to input images of the model. “Overlay” means we project the recovered mesh back onto the original images. “Cam” and “World” show the output in camera and world coordinates, respectively.
@@
Figure 8: Qualitative comparison with SPEC on SPEC-MTP (Kocabas et al., 2021b). Here are some challenging examples. Although imperfect, W-HMR shows better mesh-image alignment and higher orientation accuracy in world space.
Camera Parameter There is no such evaluation for inferred camera parameters because W-HMR cannot and does not need to predict true intrinsic and extrinsic parameters. First, focal length prediction is an ill-posed problem and lacks much training data. We choose to leverage mesh-image alignment information to predict pseudo focal lengths. This pseudo focal length may differ from the true focal length but can accurately project the recovered mesh to the human area. After all, the purpose of predicting the focal length is to handle distorted images so that the calculation of 2D joint loss is more accurate, thereby improving model performance. In summary, inferred camera parameters cannot be evaluated quantitatively because they may differ significantly from ground truth. Of course, it can be implicated by improved recovery accuracy and aligned-well qualitative results. Moreover, the module used to predict the camera rotation matrix in W-HMR comes from CamClib in SPEC. Its accuracy has been evaluated in SPEC. The novelty in our work is how we better use this information. With the same CamClib, W-HMR still significantly outperforms SPEC, which is a solid evidence to prove the effectiveness of our method.
4.4 Ablation Study
@@
Camera Calibration We select PyMAF(Zhang et al., 2021a) as our baseline and consistently use ResNet50 as the backbone throughout our ablation study, with 431 sparse vertices as the downsampling points. The training and evaluation settings are all kept the same. From the Tab. 5, it is evident that the most significant improvements stem from adding the vertex regressor and global information. Including the vertex regressor enhances the model’s ability to reconstruct body mesh details. Moreover, global information contributes independently and aids in camera calibration. A comparison between “V” and “V+F” reveals that solely adding the focal length predictor only has a noticeable impact on vertex accuracy. However, comparing “V+G” and “V+F+G” demonstrates that the presence of global information leads to more substantial improvements when combined with the focal length predictor. This finding further supports our idea that global information is crucial for recovering body orientation, predicting focal lengths, and correcting distortions. In summary, with the camera calibration assistance, W-HMR achieves a decent reconstruction accuracy, which provides a solid foundation for human body recovery in world space.
Table 5: The ablation study on the core block of the W-HMR. The experiments below utilize ResNet50 (He et al., 2016) as the backbone. In this table, “V” refers to the vertex regressor. “F” indicates whether the focal length predictor and FULL2 2D supervision are implemented. “G” signifies the addition of global information during input.
@@
Orientation Correction We select two methods for evaluation. The first method is similar to SPEC. SPEC concatenates features and the camera parameters obtained from CamCalib, then inputs them into a regressor to directly recover the human body in the world coordinate. As demonstrated in Tab. 6, the spec-like method significantly outperforms previous methods. However, we attribute this performance to overfitting. The model relies on massive parameters to memorize the global body orientation in the training set of SPEC-SYN, subsequently achieving excellent results on the test set. This is supported by the inferior performance of spec-like methods on SPEC-MTP. This indicates that such a method is prone to overfitting, leading to the misconception that the model can capture environmental information. Worse, this design destroys the model’s original learned knowledge, making the overall reconstruction less precise. In contrast, our method offers a more straightforward structure, requires less training data, and exhibits greater robustness. Furthermore, other models based on the camera coordinate can easily adopt our approach to achieve global pose recovery.
Table 6: The ablation study on body orientation correction. “spec-like” refers to using a similar method like SPEC to predict global pose directly. The numbers in brackets are the evaluation results on SPEC-MTP. This is the reason why we do not choose the spec-like method. Its impressive results are based on overfitting and sacrificing generalization.
@@
GraphCMR 181.7/181.5 86.6 219.8/218.3
SPIN 165.8/161.4 79.5 194.1/188.0
PartialHumans 169.3/174.1 88.2 207.6/210.4
I2L-MeshNet 169.8/163.3 82.0 203.2/195.9
HMR 128.7/96.4 55.9 144.2/111.8
PyMAF 126.8/88.0 48.7 136.7/118.0
@@
To verify the portability of OrientCorrect, we do related experiments, and the results are displayed in Tab. 7. To make the experiments more convincing, we choose HMR, a widely adopted model, as the baseline. We design two porting methods. One is to directly use the pre-trained OrientCorrect from W-HMR to predict the body rotation in the world space and then combine it with the pose and shape output from HMR to generate the final human motion. It can be seen that it brings nearly 14% improvement. From Tab. 2, it can be seen that using the camera rotations output from SPEC’s CamCalib to correct brings only about 10% boost.
Moreover, the most impressive is the second porting me-thod. We utilize the features and body orientation output by HMR to replace the original input part of OrientCorrect. In other words, we train an OrientCorrect adapted to HMR. The effect is pronounced, and the reconstruction accuracy is again improved by one grade. To avoid misunderstandings, you may think OrientCorrect is not as good as CamCalib or SPEC because the results in Tab. 7 are not as good as those of HMR and SPEC in Tab. 2. This is because we are based on different baselines. We use the pre-trained weights from SPIN directly and do not do any tuning, so the final results should be worse.
Table 7: Ablation study results on SPEC-MTP. The weights are from SPIN and are not fine-tuned. indicates using orientation output by OrientCorrect of pre-trained W-HMR. indicates HMR with an adaptive OrientCorrect module.
@@
Design Details To find the optimal design for our model, we conduct further ablation experiments, as shown in Tab. 8. Firstly, we test different downsampling methods. Using markers for downsampling leads to better learning of human body features. Here, “markers” refer to mocap markers. The improved performance might be because mocap markers are a better proxy for inferring human body information (Zhang et al., 2021b). However, since using mocap markers conflicts with the vertex regressor, we only employ markers in the pure parameter regression.
Table 8: Ablation experiments regarding the design details of modules. The first comparison involves vertex selection for downsampling. The second comparison examines the impact of different information inputs. The third comparison evaluates different designs of the vert graphomer. The fourth comparison evaluates different combinations of regressors. Within each comparison, all settings remain consistent except for the variations in module design.
@@
We also compare our 5D global information with the 3D bounding box information proposed by CLIFF (Li et al., 2022b). It is clear that, apart from the relative position information provided by the bounding box, the original image size also contains essential information for our task.
@@
Lastly, we test various combinations of the hybrid regression module. Both approaches show advantages, but it is apparent that the combination of “2 param$+$1 vert” has clear benefits regarding model size and inference efficiency. Therefore, we ultimately select this combination as our final regression module.
@@
5 Discussion
Difference It is not the first time that predicting the distance between a person and a camera has been proposed. To understand our innovation, we take (Moon et al., 2019) as an example to discuss the differences between our work. Different Methods: The neural network in our work directly predicts the person’s distance from the camera. But they first calculate values $k$ based on the sizes of the bounding box and image. Then, they predict a correction factor $\gamma$ through a neural network and get the absolute depth value using $k$ and $\gamma$. Different Theories: Our work mainly relies on the degree of distortion of humans in the images. The model tends to predict shorter distances when humans are more distorted. But (Moon et al., 2019) mainly relies on the human size in images. A more miniature human means this human stands farther away from the camera. Different Purposes: We use predicted distance to calculate focal lengths. Our ultimate goal is to achieve a perspective camera model so that we can make full use of distorted images. As for (Moon et al., 2019), they want to recover accurate relative positions for multi-person pose estimation.